OpenStudy (anonymous):

Find the perimeter of this trapezoid. A(2, 1), B(6, 5), C(3, 5), D(1, 3)

OpenStudy (mertsj):

Do you know the distance formula?

Parth (parthkohli):

Use the distance formula for all the vertices and add them up.

OpenStudy (anonymous):

no

OpenStudy (mertsj):

I will demonstrate one time and then you can practice.

OpenStudy (anonymous):

I mean, I don't know how to use it for this..

OpenStudy (anonymous):

@Mertsj thanks

Parth (parthkohli):

\(\Large \color{MidnightBlue}{\Rightarrow D = \sqrt{(x_1 - x_2)^2 + (y_1 - y_2)^2} }\)

OpenStudy (anonymous):

what about perimeter?

OpenStudy (anonymous):

is that going to be the same as diameter for this?

Parth (parthkohli):

You add up the distances of all the lines.

OpenStudy (mertsj):

Choose two of the given points and plug into the distance formula.

OpenStudy (mertsj):

Start with A and B

Parth (parthkohli):

Sketch it, then see all the vertices. Apply the distance formula for all lines and then add the distances to get the perimeter.

OpenStudy (mertsj):

That will give you the length of side AB.

Parth (parthkohli):

@Mertsj is a perfect teacher, I'll let him continue

OpenStudy (anonymous):

I think he left!

OpenStudy (anonymous):

alright.. sketching..

OpenStudy (mertsj):

\[AB=\sqrt{(6-2)^2+(5-1)^2}=\sqrt{16+16}=4\sqrt{2}\]

Parth (parthkohli):

@rokotherodent did you understand?

OpenStudy (mertsj):

Now you try BC

OpenStudy (anonymous):

A little bit..

OpenStudy (mertsj):

And enough of this talk about diameter. Diameter is a characteristic of a circle. We're dealing with a trapezoid here.

OpenStudy (anonymous):

@ParthKohli I'm a little confused..

OpenStudy (anonymous):

Thanks @Mertsj because I wanted perimieter

Parth (parthkohli):

I may help you understand. Where are you confused btw?

OpenStudy (anonymous):

well, I've drawn the trapezoid.. but then

OpenStudy (mertsj):

So find the length of BC. It's an easy one because 5-5=0

Parth (parthkohli):

Think about the distance between two points by using the distance formula.

OpenStudy (anonymous):

what do I need? Do I need AB, BC, CD, and thats all?

Parth (parthkohli):

When you add the lengths of all sides, you get the perimeter.

OpenStudy (mertsj):

Remember that perimeter is the distance around. So find the length of each side and add them together.

Parth (parthkohli):

Trapezoid = four sides

OpenStudy (anonymous):

so BC = sqrt 9 (3)?

OpenStudy (mertsj):

\[BC=\sqrt{(6-3)^2+(5-5)^2}=\sqrt{3^2+0^2}=\sqrt{9}=3\]

Parth (parthkohli):

(6,5)(3,5) \(\Large \color{MidnightBlue}{\Rightarrow \sqrt{(6 - 3)^2 + 0^2} }\) \(\Large \color{MidnightBlue}{\Rightarrow \sqrt{3^2} }\) \(\Large \color{MidnightBlue}{\Rightarrow 3 }\) Am I correct?

OpenStudy (anonymous):

Yes, you both are!

OpenStudy (anonymous):

This is what I have..

Parth (parthkohli):

Can you do it now?

OpenStudy (anonymous):

Now lastly, I have one more question: the blank has 6 sqrt _

OpenStudy (anonymous):

do I just put 3 in that blank then?

Parth (parthkohli):

You add the lengths of all sides. Remember.

OpenStudy (mertsj):

So far we have this start: Perimeter = AB + BC + CD + AD = 4sqrt2+3 + CD + AD

OpenStudy (mertsj):

So if we could just find CD and AD we would have it!!

OpenStudy (anonymous):

ok

Parth (parthkohli):

Yep!

OpenStudy (anonymous):

CD is sqrt 4

OpenStudy (anonymous):

AD is sqrt 2

Parth (parthkohli):

Hmm no

OpenStudy (mertsj):

\[CD=\sqrt{(5-3)^2+(3-1)^2}=\sqrt{2^2+2^2}=\sqrt{4+4}=\sqrt{8}=2\sqrt{2}\]

OpenStudy (mertsj):

So now we have Perimeter = AB + BC+CD+AD=4sqrt2+3+2sqrt2+AD Try AD again.

OpenStudy (mertsj):

Follow the distance formula.

OpenStudy (anonymous):

ok 2 - 1 = 1 1 - 3 = neg. 2

Parth (parthkohli):

@Mertsj I'm 99 now ^_^

OpenStudy (mertsj):

Squaare them now.

OpenStudy (anonymous):

okay, 2 squared is 4

Parth (parthkohli):

And square them....then find the square root

OpenStudy (mertsj):

What is 1 squared?

Parth (parthkohli):

What is \(\sqrt{1 + 4}\)?

OpenStudy (anonymous):

1 + 4 = 5

Parth (parthkohli):

Yes. Now add all the lengths.

OpenStudy (mertsj):

Good. So AD=?

OpenStudy (anonymous):

sqrt 5?

Parth (parthkohli):

Yes

OpenStudy (anonymous):

AD = sqrt 5

OpenStudy (anonymous):

now, adding them together..

OpenStudy (mertsj):

Yes. Now add all four sides and remember that only similar radicals can be added.

Parth (parthkohli):

Now add the side lengths: 4sqrt2+3+2sqrt2+sqrt5

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

so that will be 6 sqrt 2 + 3 sqrt 5?

OpenStudy (mertsj):

@ParthKohli I see that you are 99 now. How did that happen so quickly?

OpenStudy (mertsj):

Well but, I don't think you should change 3 + sqrt5 to 3 TIMES sqrt5, do you?

Parth (parthkohli):

Result of hard work from the day I joined the site aka 22 April lol

OpenStudy (anonymous):

ok I agree @Mertsj

OpenStudy (anonymous):

I think that the final answer will be: 6 sqrt 2 + sqrt 5 + 3

OpenStudy (mertsj):

So why don't we just leave it 6sqrt2+3+sqrt5 since there are no other like terms that can be added.

OpenStudy (anonymous):

Well, I thought that we could in fact leave it at that.

OpenStudy (mertsj):

And you are correct and now we know how far it is around that trapezoid.!!!

Parth (parthkohli):

Yes, you could. Unlike terms.

Parth (parthkohli):

Finally done...ahh

OpenStudy (anonymous):

Thanks to both! I appreciate your help!

Parth (parthkohli):

you're welcome :D

OpenStudy (mertsj):

yw and we hope you have learned a lot.

Parth (parthkohli):

Can you do the rest?

OpenStudy (anonymous):

I have! I have an identical problem that I'll work myself! :D @ParthKohli

Parth (parthkohli):

Sure....! If you have any confusion then contact me or Mertsj.

OpenStudy (mertsj):

Excellent. And then if you want someone to check it you have ParthKohli's number.

OpenStudy (anonymous):

Alrighty! thanks!